OpenStudy (anonymous):

What is the solution set for |–x + 2| = 5

OpenStudy (anonymous):

|–x + 2| = 5 -x+2=5 , -x+2=-5 -x=3 , -x=-7 x=-3 , x=7

OpenStudy (anonymous):

when u want to get ride of a modulus ,put \[\pm \] instead